# FlexiClip: Locality-Preserving Free-Form Character Animation

### v1

@@ Abstract
Animating clipart images with seamless motion while maintaining visual fidelity and temporal coherence presents significant challenges. Existing methods, such as AniClipart, effectively model spatial deformations but often fail to ensure smooth temporal transitions, resulting in artifacts like abrupt motions and geometric distortions. Similarly, text-to-video (T2V) and image-to-video (I2V) models struggle to handle clipart due to the mismatch in statistical properties between natural video and clipart styles. This paper introduces FlexiClip, a novel approach designed to overcome these limitations by addressing the intertwined challenges of temporal consistency and geometric integrity. FlexiClip extends traditional Bézier curve-based trajectory modeling with key innovations: temporal Jacobians to correct motion dynamics incrementally, continuous-time modeling via probability flow ODEs (pfODEs) to mitigate temporal noise, and a flow matching loss inspired by GFlowNet principles to optimize smooth motion transitions. These enhancements ensure coherent animations across complex scenarios involving rapid movements and non-rigid deformations. Extensive experiments validate the effectiveness of FlexiClip in generating animations that are not only smooth and natural but also structurally consistent across diverse clipart types, including humans and animals. By integrating spatial and temporal modeling with pre-trained video diffusion models, FlexiClip sets a new standard for high-quality clipart animation, offering robust performance across a wide range of visual content. Project Page: https://flexiclip.github.io/
@@ 1 Introduction
To overcome these limitations, we introduce FlexiClip, a novel approach that extends the state of the art by addressing the key challenges of temporal coherence and geometric consistency in animated clipart. FlexiClip also builds on the basic framework of modeling keypoint trajectories with cubic Bézier curves but introduces significant innovations to improve temporal dynamics and maintain a consistent animation pipeline. Central to FlexiClip is the use of temporal Jacobian for incremental temporal corrections, pfODE (Song et al., 2020; de Albuquerque & Pearson, 2024) for continuous-time integration of these corrections, and a flow matching loss inspired by GFlowNet (Bengio et al., 2023) to ensure smooth temporal evolution and reduction of temporal noise. Although AniClipart also represents keypoint motion through Bézier curves and employs ARAP (as rigid as possible) deformation to model spatial consistency, its temporal modeling lacks a mechanism for addressing noise accumulation across frames. In contrast, FlexiClip introduces a novel paradigm by learning temporal Jacobians that incrementally correct the spatial Jacobian over time. This framework enables precise control over temporal evolution and prevents drift in the motion dynamics, thereby maintaining the animation’s naturalness and consistency across longer sequences. Additionally, we leverage probability flow ODE (pfODE) (Sec.2.2) to model the temporal correction process as a continuous-time function, addressing temporal noise more effectively than discrete-time optimization methods.
@@
A critical challenge in temporal animation is the preservation of motion smoothness without introducing geometric distortions. In previous works, such as AniClipart, motion dynamics are modeled independently for each frame, which can lead to inconsistent transitions and visual artifacts (Fig 3). In contrast, FlexiClip utilizes pfODE (Sec.2.2) to model the evolution of temporal Jacobian, which are corrected progressively over time. This continuous correction mechanism ensures that the temporal noise is mitigated, resulting in smoother and more consistent motion transitions. Moreover, this framework provides a novel solution to the problem of maintaining structural consistency (locality preserving) during fast spatial transitions, which is often a difficult task in motion modeling. Another novel contribution of FlexiClip is the flow matching loss, which leverages the principles of GFlowNet (Bengio et al., 2023) to optimize the temporal noise reduction process. The flow matching loss operates by comparing the dynamics of the spatial Jacobian and their temporal corrections over time. This approach ensures that temporal noise introduced by the spatial model is progressively reduced, facilitating smoother transitions between frames and preventing the accumulation of errors. Unlike earlier approaches, such as AniClipart, that do not explicitly model the evolution of temporal noise, FlexiClip provides a robust mechanism for controlling noise accumulation and preserving the underlying motion’s coherence. At last, FlexiClip integrates learning of spatial and temporal modules with the concept of video Score Distillation Sampling (SDS), which allows us to distill the knowledge of pre-trained video diffusion models to guide the animation generation process. Extensive experiments and ablation studies demonstrate FlexiClip’s ability to generate smooth, natural, and temporally consistent clipart animations across a wide range of visual content, including humans and animals (see Fig. 1). FlexiClip also supports handling complex animations (Fig. 5) where keypoint transitions involve non-rigid deformations like rotation, complex motion, etc. In summary, our contributions can be summarized as follows:
@@ 2.2 Probability flow ODE (pfODE)
where $f(x,t)$ is the drift term, $G(x,t)$ represents the noise diffusion matrix, and $\nabla_{x}\log p_{t}(x)$ is the score function, approximated by the diffusion model. This reverse process reconstructs the data distribution by denoising, guided by the score function $\nabla_{x}\log p_{t}(x)$. However, in addition to the SDE (Song et al., 2020) proposed the probability flow ODE (pfODE), which satisfies the same Fokker-Planck equation, but is deterministic (no noise term). it is given by:
$\frac{dx}{dt}=\left(f(x,t)-\frac{1}{2}\nabla\cdot\left[G(x,t)G(x,t)^{T}\right]\right.$
$\left.-\frac{1}{2}G(x,t)G(x,t)^{T}\nabla_{x}\log p_{t}(x)\right)$ (8)
This pfODE evolves the data points smoothly without the stochastic Brownian motion term. This process is deterministic, and data points evolve smoothly, resulting in a flow that preserves local neighborhoods. Under the Gaussian noise assumption, the score function $\nabla_{x}\log p_{t}(x)$ can be trained from a pre-trained diffusion model via score matching, and hence can guide mesh deformation starting with noisy vertices. However, the key challenge in learning the mesh deformation directly with pre-trained diffusion models is that these models reverse the isotropic Gaussian noise, as described in the SDE-based formulation, making it harder to directly match the target distribution (highly structured, like posed mesh). It is straightforward to show (App. C) that the class of time-varying densities satisfies (8) when $f=0$ and $GG^{T}=\dot{C}$, similar to (de Albuquerque & Pearson, 2024) we consider the pfODE with rescaling to avoid variance overflow:
@@ 2.3 GFlowNets
Generative Flow Networks (GFlowNets) (Bengio et al., 2023) enable training generative models with unnormalized target densities. A GFlowNet is represented as a directed acyclic graph $G=(S,A)$, where $S$ is the set of states and $A\subseteq S\times S$ is the set of actions. Transitions between states are deterministic, with an initial state $s_{0}$ and terminal states $s_{N}$. The forward policy $P_{F}(s^{\prime}|s)$ defines the transition from $s$ to $s^{\prime}$, while the backward policy $P_{B}(s|s^{\prime})$ defines the reverse. The goal is to learn a forward policy such that the terminal state distribution $P_{T}(x)\propto R(x)$, where $R(x)$ is the unnormalized reward function.
@@ 3 FlexiClip
In this section, we introduce our mesh deformation framework, FlexiClip (Fig.2). We begin with a method overview (Sec. 3.1), followed by a spatial posing with Jacobian Fields on along parameterized Bézier trajectories (Sec. 3.2). Next, we discuss how temporal signals are modeled using pfODE to ensure temporally coherent motion (Sec. 3.3). Finally, we outline the loss functions used (Sec. 3.4).
@@ 3.3 Temporal Smoothing
ODE Formulation: Since the first frame is stationary and given as input, we set the first frame’s temporal Jacobian as, $\mathbf{J}_{0}^{R}=\mathbf{0}\in\mathbb{R}^{F\times 2\times 2}$. Furthermore, we are required to remove only temporal noise from spatial Jacobians and wanted to add only temporal correction term (temporal Jacobian), we propose to reformulate the ODE given in (Eq.9) given as:
@@
Here, $\mathbf{J}_{0}^{R}$ is initialized to zero for the first frame, ensuring a neutral starting state. Numerical methods such as Euler’s method are used for integration. In comparison to (Eq.9) we model the time-varying noise $C(t)$ with $C_{W}^{P}$ and $C_{W-1}^{R}$. $C_{W}^{P}$ is the attention over spatial Jacobians over a time window $W$ i.e. till the current time step. And, $C_{W-1}^{R}$ is the attention over temporal Jacobian in a time window $W-1$ i.e. till previous step. From (Eq.9) we need score function which can estimate a de-noised version of a noise-corrupted data sample given noise level $C(t)$, similarly (Eq.12) denoise the input noise accumulated over spatial Jacobians $C_{W}^{P}\rightarrow C(t)$ and the corrections applied to them $C_{W-1}^{R}\rightarrow\textrm{rescaling }A(t)$ we are estimating the denoised Jacobian which is the correction term for spatial Jacobian. Hence, as in (Eq.9) as soon as the corrections applied starts to contract it balance the opposing force from the error in spatial Jacobians and finally the corrections term reduced to zero. To ensure this correction term reduces to zero only for temporal noise we apply the flow matching (detailed balance) objective from GFlowNets, detailed subsequently.
@@ 3.4 Loss function
We distill the prior knowledge of a pretrained video diffusion model via Video Score Distillation Sampling (SDS). Let $\mathbf{J}=\{\mathbf{J}_{t}\}_{t=0}^{N-1}$ (Eq.10) represent the Jacobian fields over time, and $\mathbf{V}^{*}=\{\mathbf{V}_{t}^{*}\}_{t=0}^{N-1}$ be the temporally consistent set of vertices computed via differentiable function $g$ given $\mathbf{J}_{t}$ and $\mathbf{T}_{c}$ for each time step (Sec.2.1). For each time step $t$, the deformed mesh $\mathbf{M}_{t}^{*}=(\mathbf{V}_{t}^{*},\mathbf{F})$ is used to warp $\mathcal{W}$ initial clipart image and then rendered using a differentiable renderer $\mathcal{R}$ (Li et al., 2020), producing frames $\mathbf{I}_{t}=\mathcal{R}(\mathcal{W}(I_{0}),\mathbf{M}_{t}^{*})$. The sequence of frames forms the video $\mathbf{X}=\{\mathbf{I}_{t}\}_{t=0}^{N-1}$. The pretrained video diffusion model ${\epsilon}_{\phi}$ with the input video $\mathbf{X}$ as parameters, it produces a gradient w.r.t $\theta$ which are the spatial and temporal parameters driving the animation:
@@
Flow Matching Loss: Optimizing the mesh deformation with Video SDS loss using Gaussian noise assumption can produce animations with abrupt transitions, exhibit local geometric distortions as the displacement for each keypoint is predicted independently and hence not be able to match the target distribution (as explained in Sec.2.2). In conclusion, we are required to optimize the temporal noise (Sec.3.3) for smooth evolution of keypoints and maintain local structure, however these pre-trained video diffusion models under isotropic Gaussian noise assumption increase the effective dimensionality of the data, which may begin as a low-dimensional manifold embedded within same dimensionality. Thus, maintaining intrinsic data dimensionality requires a choice of flow that preserves this dimension. We define the dimensionality as the denoising objective, for example, consider the noise quantified as $C(t)$
@@
where $D(\cdot)$ de-noised version of a noise-corrupted data sample $\mathbf{X}$ given noise level $C(t)$. In practice we used the same denoiser $\epsilon_{\phi}$ U-Net of pre-trained video diffusion model, hence (Eq.15) is essentially the score function $\nabla_{\mathbf{X}}\log p(\mathbf{X})$. To preserve the dimension for the temporal noise, despite learning the deformation under the Gaussian noise assumption, we propose to take the difference between the score function obtained for frames $\mathbf{I}_{t}$ produced through overall Jacobian $\mathbf{J}_{t}$ and spatial Jacobian $\mathbf{J}_{t}^{P}$. This loss is inspired from the detailed balance (DB) objective from GFlowNets where, the backward process (Spatial Jacobian) keeps on introducing the temporal noise and the forward process keeps on reducing the added temporal noise for each state transition. Since the score function is already normalized and comparable there is no normalizing factor $F(s)$ as in Sec.2.3, thus the loss function is given as:
@@
The Overall Loss for FlexiClip is defined as the weighted sum loss in Eq.13, 15, 16:

### v2

@@ Abstract
Animating clipart images with seamless motion while maintaining visual fidelity and temporal coherence presents significant challenges. Existing methods, such as AniClipart, effectively model spatial deformations but often fail to ensure smooth temporal transitions, resulting in artifacts like abrupt motions and geometric distortions. Similarly, text-to-video (T2V) and image-to-video (I2V) models struggle to handle clipart due to the mismatch in statistical properties between natural video and clipart styles. This paper introduces FlexiClip, a novel approach designed to overcome these limitations by addressing the intertwined challenges of temporal consistency and geometric integrity. FlexiClip extends traditional Bézier curve-based trajectory modeling with key innovations: temporal Jacobians to correct motion dynamics incrementally, continuous-time modeling via probability flow ODEs (pfODEs) to mitigate temporal noise, and a flow matching loss inspired by GFlowNet principles to optimize smooth motion transitions. These enhancements ensure coherent animations across complex scenarios involving rapid movements and non-rigid deformations. Extensive experiments validate the effectiveness of FlexiClip in generating animations that are not only smooth and natural but also structurally consistent across diverse clipart types, including humans and animals. By integrating spatial and temporal modeling with pre-trained video diffusion models, FlexiClip sets a new standard for high-quality clipart animation, offering robust performance across a wide range of visual content. Project Page: https://creative-gen.github.io/flexiclip.github.io/
@@ 1 Introduction
To overcome these limitations, we introduce FlexiClip, a novel approach that extends the state of the art by addressing the key challenges of temporal coherence and geometric consistency in animated clipart. FlexiClip also builds on the basic framework of modeling keypoint trajectories with cubic Bézier curves but introduces significant innovations to improve temporal dynamics and maintain a consistent animation pipeline. Central to FlexiClip is the use of temporal Jacobian for incremental temporal corrections, pfODE (Lim et al., 2023; de Albuquerque & Pearson, 2024) for continuous-time integration of these corrections, and a flow matching loss inspired by GFlowNet (Bengio et al., 2023) to ensure smooth temporal evolution and reduction of temporal noise.
Although AniClipart also represents keypoint motion through Bézier curves but employs ARAP (As Rigid As Possible) deformation to model spatial consistency, its temporal modeling lacks a mechanism for addressing noise accumulation across frames. In contrast, FlexiClip introduces a novel paradigm by learning temporal Jacobians that incrementally correct the spatial Jacobian over time. This framework enables precise control over temporal evolution and prevents drift in the motion dynamics, thereby maintaining the animation’s naturalness and consistency across longer sequences. Additionally, we leverage probability flow ODE (pfODE) (Sec.2.2) to model the temporal correction process as a continuous-time function, addressing temporal noise more effectively than discrete-time optimization methods.
@@
A critical challenge in temporal animation is the preservation of motion smoothness without introducing geometric distortions. In previous works, such as AniClipart, motion dynamics are modeled independently for each frame, which can lead to inconsistent transitions and visual artifacts (Fig 3). In contrast, FlexiClip utilizes pfODE (Sec.2.2) to model the evolution of temporal Jacobian, which are corrected progressively over time. This continuous correction mechanism ensures that the temporal noise is mitigated, resulting in smoother and more consistent motion transitions. Moreover, this framework provides a novel solution to the problem of maintaining structural consistency (locality preserving) during fast spatial transitions, which is often a difficult task in motion modeling.
Another novel contribution of FlexiClip is the flow matching loss, which leverages the principles of GFlowNet (Bengio et al., 2023) to optimize the temporal noise reduction process. The flow matching loss operates by comparing the dynamics of the spatial Jacobian and their temporal corrections over time. This approach ensures that temporal noise introduced by the spatial model is progressively reduced, facilitating smoother transitions between frames and preventing the accumulation of errors. Unlike earlier approaches, such as AniClipart, that do not explicitly model the evolution of temporal noise, FlexiClip provides a robust mechanism for controlling noise accumulation and preserving the underlying motion’s coherence.
At last, FlexiClip integrates learning of spatial and temporal modules with the concept of video Score Distillation Sampling (SDS), which allows us to distill the knowledge of pre-trained video diffusion models to guide the animation generation process. Extensive experiments and ablation studies demonstrate FlexiClip’s ability to generate smooth, natural, and temporally consistent clipart animations across a wide range of visual content, including humans and animals (see Fig. 1). FlexiClip also supports handling complex animations (Fig. 5) where keypoint transitions involve non-rigid deformations like rotation, complex motion, etc. In summary, our contributions can be summarized as follows:
@@ 2.1 Representing Shapes as Jacobian Fields
For a triangular face $f$ with vertices $\{\mathbf{v}_{i},\mathbf{v}_{j},\mathbf{v}_{k}\}$, the Jacobian matrix captures how the vertex positions within the triangle are influenced by barycentric coordinates:
$\mathbf{J}^{*}_{f}=\begin{bmatrix}\frac{\partial\mathbf{v}_{i}^{*}}{\partial\mathbf{v}_{i}}&\frac{\partial\mathbf{v}_{j}^{*}}{\partial\mathbf{v}_{i}}&\frac{\partial\mathbf{v}_{k}^{*}}{\partial\mathbf{v}_{i}}\\ \frac{\partial\mathbf{v}_{i}^{*}}{\partial\mathbf{v}_{j}}&\frac{\partial\mathbf{v}_{j}^{*}}{\partial\mathbf{v}_{j}}&\frac{\partial\mathbf{v}_{k}^{*}}{\partial\mathbf{v}_{j}}\\ \frac{\partial\mathbf{v}_{i}^{*}}{\partial\mathbf{v}_{k}}&\frac{\partial\mathbf{v}_{j}^{*}}{\partial\mathbf{v}_{k}}&\frac{\partial\mathbf{v}_{k}^{*}}{\partial\mathbf{v}_{k}}\end{bmatrix}.$ (6)
Updated Jacobians give local deformation of each triangle.
@@ 2.2 Probability flow ODE (pfODE)
where $f(x,t)$ is the drift term, $G(x,t)$ represents the noise diffusion matrix, $\bar{W}$ is time-reversed Brownian motion and $\nabla_{x}\log p_{t}(x)$ is the score function, approximated by the diffusion model. This reverse process reconstructs the data distribution by denoising, guided by the score function $\nabla_{x}\log p_{t}(x)$. However, in addition to the SDE, Song et al. (Lim et al., 2023) proposed the probability flow ODE (pfODE), which satisfies the same Fokker-Planck equation, but is deterministic (no Brownian term). it is given by:
$\frac{dx}{dt}=\left(f(x,t)-\frac{1}{2}\nabla\cdot\left[G(x,t)G(x,t)^{T}\right]\right.\\ \left.-\frac{1}{2}G(x,t)G(x,t)^{T}\nabla_{x}\log p_{t}(x)\right)$ (9)
This pfODE evolves the data points smoothly without the stochastic Brownian motion term. This process is deterministic, and data points evolve smoothly, resulting in a flow that preserves local neighborhoods.
Under the Gaussian noise assumption, the score function $\nabla_{x}\log p_{t}(x)$ can be trained from a pre-trained diffusion model via score matching, and hence can guide mesh deformation starting with noisy vertices. However, the key challenge in learning the mesh deformation directly with pre-trained diffusion models is that these models reverse the isotropic Gaussian noise, as described in the SDE-based formulation, making it harder to directly match the target distribution (highly structured, like posed mesh). It is straightforward to show (App. C) that the class of time-varying densities satisfies (9) when $f=0$ and $GG^{T}=\dot{C}$, similar to (de Albuquerque & Pearson, 2024) we consider the pfODE with rescaling to avoid variance overflow:
@@ 2.3 GFlowNets
Generative Flow Networks (GFlowNets) (Bengio et al., 2023) enable training generative models with unnormalized target densities. GFlowNet is represented as a directed acyclic graph $G=(S,A)$, where $S$ is the set of states and $A\subseteq S\times S$ is the set of actions. Transitions between states are deterministic, with an initial state $s_{0}$ and terminal states $s_{N}$. The forward policy $P_{F}(s^{\prime}|s)$ defines the transition from $s$ to $s^{\prime}$, while the backward policy $P_{B}(s|s^{\prime})$ defines the reverse. The goal is to learn a forward policy such that the terminal state distribution $P_{T}(x)\propto R(x)$, where $R(x)$ is the unnormalized reward function.
@@
Generative Flow Networks (GFlowNets) (Bengio et al., 2023) provide a framework for training generative models with an unnormalized target density function. A GFlowNet is represented as a directed acyclic graph $G=(S,A)$, where $S$ is the set of states and $A\subseteq S\times S$ is the set of actions. The transition between states is deterministic, and the network has an initial state $s_{0}$ and terminal states $s_{N}$. The forward policy $P_{F}(s^{\prime}|s)$ defines the transition probability from state $s$ to $s^{\prime}$, while the backward policy $P_{B}(s|s^{\prime})$ defines the reverse transition. The goal is to learn a forward policy such that the terminal state distribution $P_{T}(x)\propto R(x)$, where $R(x)$ is an unnormalized reward function.
The flow function $F(\tau)$ incorporates the normalizing factor, and the state flow function $F(s)$ models the unnormalized probability flow for each state. GFlowNet’s training uses the detailed balance (DB) condition, which ensures that for any transition $(s\to s^{\prime})$, the following holds: $F(s)P_{F}(s^{\prime}|s)=F(s^{\prime})P_{B}(s|s^{\prime}),\quad\forall(s\to s^{\prime})\in A.$ For terminal states $x$, the condition $F(x)=R(x)$ is required. Satisfying this DB criterion ensures that the terminal distribution $P_{T}(x)$ matches the desired target distribution, proportional to $R(x)$.
@@ 3 FlexiClip
In this section, we introduce our mesh deformation framework, FlexiClip (Fig.2). We begin with a method overview (Sec. 3.1), followed by a spatial posing with Jacobian Fields on parameterized Bézier trajectories (Sec. 3.2). Next, we discuss how temporal signals are modeled using pfODE to ensure temporally coherent motion (Sec. 3.3). Finally, we outline the loss functions used (Sec. 3.4).
@@ 3.3 Temporal Smoothing
ODE Formulation:
Since the first frame is stationary and given as input, we set the first frame’s temporal Jacobian as, $\mathbf{J}_{0}^{R}=\mathbf{0}\in\mathbb{R}^{F\times 2\times 2}$. Furthermore, we are required to remove only temporal noise from spatial Jacobians and wanted to add only temporal correction term (temporal Jacobian), we propose to reformulate the ODE given in (Eq.10) given as:
@@
Here, $\mathbf{J}_{0}^{R}$ is initialized to zero for the first frame, ensuring a neutral starting state. Numerical methods such as Euler’s method are used for integration. In comparison to (Eq.10) we model the time-varying noise $C(t)$ with $C_{W}^{P}$ and rescaling $A(t)$ with $C_{W-1}^{R}$. $C_{W}^{P}$ is the attention over spatial Jacobians over a time window $W$ i.e. till the current time step. And, $C_{W-1}^{R}$ is the attention over temporal Jacobian in a time window $W-1$ i.e. till previous step. In (Eq.10) we need score function which can estimate a de-noised version of a noise-corrupted data sample given noise level $C(t)$. Comparatively, (Eq.13) also denoise the input noise accumulated over spatial Jacobians $C_{W}^{P}\rightarrow C(t)$ and the corrections applied to them $C_{W-1}^{R}\rightarrow\textrm{rescaling }A(t)$ we are estimating the denoised Jacobian which is the correction term for spatial Jacobian. For convergence, as in (Eq.10) as soon as the corrections applied starts to contract it balance the opposing force from the error in spatial Jacobians and finally the corrections term reduced to zero. To ensure this correction term reduces to zero only for temporal noise we apply the flow matching (detailed balance) objective from GFlowNets, detailed subsequently.
@@ 3.4 Loss function
We distill the prior knowledge of a pretrained video diffusion model via Video Score Distillation Sampling (SDS). Let $\mathbf{J}=\{\mathbf{J}_{t}\}_{t=0}^{N-1}$ (Eq.11) represent the predicted Jacobian fields over time, and $\mathbf{V}^{*}=\{\mathbf{V}_{t}^{*}\}_{t=0}^{N-1}$ be the temporally consistent set of vertices computed via differentiable function $g$ given $\mathbf{J}_{t}$ and $\mathbf{T}_{c}$ for each time step (Sec.2.1). For each time step $t$, the deformed mesh $\mathbf{M}_{t}^{*}=(\mathbf{V}_{t}^{*},\mathbf{F})$ is used to warp $\mathcal{W}$ initial clipart image and then rendered using a differentiable renderer $\mathcal{R}$ (Li et al., 2020), producing frames $\mathbf{I}_{t}=\mathcal{R}(\mathcal{W}(I_{0}),\mathbf{M}_{t}^{*})$. The sequence of frames forms the video $\mathbf{X}=\{\mathbf{I}_{t}\}_{t=0}^{N-1}$. The pretrained video diffusion model ${\epsilon}_{\phi}$ with the input video $\mathbf{X}$ as parameters, it produces a gradient w.r.t $\theta$ which are the spatial and temporal parameters driving the animation:
@@
Flow Matching Loss
Optimizing the mesh deformation with Video SDS loss using Gaussian noise assumption can produce animations with abrupt transitions, exhibit local geometric distortions as the displacement for each keypoint is predicted independently and hence not be able to match the target distribution (as explained in Sec.2.2). In conclusion, we are required to optimize the temporal noise (Sec.3.3) for smooth evolution of keypoints and maintain local structure, however these pre-trained video diffusion models under isotropic Gaussian noise assumption increase the effective dimensionality of the data, which may begin as a low-dimensional manifold embedded within same dimensionality. Thus, maintaining intrinsic data dimensionality requires a choice of flow that preserves this dimension. We define the dimensionality as the denoising objective, for example, consider the noise quantified as $C(t)$
@@
where $D(\cdot)$ de-noised version of a noise-corrupted data sample $\mathbf{X}$ given noise level $C(t)$. In practice we used the same denoiser $\epsilon_{\phi}$ U-Net of pre-trained video diffusion model, hence (Eq.16) is essentially the score function $\nabla_{\mathbf{X}}\log p(\mathbf{X})$.
To preserve the dimension for the temporal noise, despite learning the deformation under the Gaussian noise assumption, we propose to take the difference between the score function obtained for frames $\mathbf{I}_{t}$ produced through overall Jacobian $\mathbf{J}_{t}$ and spatial Jacobian $\mathbf{J}_{t}^{P}$. This loss is inspired from the detailed balance (DB) objective from GFlowNets where, the backward process (Spatial Jacobian) keeps on introducing the temporal noise and the forward process keeps on reducing the added temporal noise for each state transition. Since the score function is already normalized and comparable there is no normalizing factor $F(s)$ as in Sec.2.3, thus the loss function is given as:
@@
The Overall Loss for FlexiClip is defined as the weighted sum loss in Eq.14, 16, 17: